OpenStudy (anonymous):

an object is thrown straight up from the top of a tall building with a speed of 10.5 m/s. how high above or how low below the projection point is the object after 2.5s?

OpenStudy (anonymous):

acceleration due to gravity=9.8m/s2 initial velocity=-10.5m/s (the negative sign means that the object is moving against gravity) and time taken is 2.5m/s With this information, can you figure out which equation to use to find the distance?

OpenStudy (sumeet):

something is wrong with your data. time taken can't be m/s...

OpenStudy (anonymous):

oh, yeah its 2.5s unintentional mistake

OpenStudy (sumeet):

First use \[v = u + at\] where v= velocity after time t t = time lapsed u =initial velocity object is thrown upwards to g=-9.81 u=10.5 m/s Calculate v, i.e. velocity after 2.5 s Then use \[v^2 = u^2 + 2as\] a=-9.81 u=10.5m/s v= as found out using above equation s=displacement in m

OpenStudy (anonymous):